The resistance corresponding to the lower and the upper fixed points, in a platinum resistance thermometer are 3.50Ω and 3.65Ω what would be the resistance at a temperature equal to the freezing point of mercury(-37 °)? 


I hav tried this quest but im little bit confused

Grade:12

1 Answers

Vikas TU
10 years ago
Assuming the fixed points are for 0ºC and 100ºC, then ..

?R ≡ 100ºC 
1ºC = (?R / 100)Ω = {(3.65 - 3.50) / 100}Ω .. .. 1ºC ≡ 0.0015Ω

37º below Ro = (37 x 0.0015)Ω below Ro = R
R = Ro - (37 x 0.0015)Ω .. .. (Ro = resistance at 0º, 3.50Ω)

R = 3.50Ω - (37 x 0.0015)Ω .. .. ?R = 3.44Ω
  • PLZ  APPROVE!!!!!!!!!!!!!!1